Find the number of sides of a regular convex polygon whose angle is 40degrees.

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Hemant

Total angle of polygon=(n-2) × 180°

each interior angle=(n-2) × 180°/n

here each interior angle=180-40=140

(n-2) × 180°/n = 140 solving we get n=9

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..